OpenStudy (anonymous):

Could someone help me with these? I thnk I'm missing a step \[(3\sqrt{x+5})+2x ^{3}-6\] \[(3\sqrt{x+5})-2x ^{3}-6\]

OpenStudy (anonymous):

well, you could start with stating the problem. Is this a system of equations?

OpenStudy (anonymous):

Functions. The origional statement is: Given \[f(x)=3\sqrt{x+5} \] and \[g(x)=2x ^{3}-6\] find (f+g)(4)

OpenStudy (anonymous):

Well, (f+g) means that you just have to add those functions. (f+g)(4) means that you have to find the value of the sum at point x=4. \[(f+g)(4)=3\sqrt{4+5}+2*4^3-6=3*3+128-6=9+122=131\]

OpenStudy (anonymous):

Oh, I see. I was misreading it, adding the equations then trying to multiply by four You put the 4 in first then add. Ok, that makes more sense

OpenStudy (anonymous):

Thank you